A chemist prepares a solution of magnesium fluoride by measuring out 0.010 mol of magnesium fluoride into a 100 mL volumetric flask and filling the flask to the mark with water.
Calculate the concentration in mol/L of the chemist's magnesium fluoride solution. Be sure your answer has the correct number of significant digits.
1
Expert's answer
2020-06-08T15:27:04-0400

V= 100 mL = 0.100 L; n = 0.010 mol

C = n/V = 0.010/0.100 = 0.10 mol/L